OpenStudy (anonymous):

Use the method of cylindrical shells to find the volume V of the solid obtained by rotating the region bounded by the given curves about the x-axis. x = 1 + (y − 6)2, x = 10

OpenStudy (amistre64):

i remember these :)

OpenStudy (anonymous):

first, you should graph the given curves and find the points of intersection

OpenStudy (anonymous):

Here's the graph

OpenStudy (anonymous):

I calculated that the intersections are at (10,3) and (10,9)

OpenStudy (amistre64):

i wonder if moving it is a good idea ... prolly not at the moment tho x = 1 + (y − 6)2, x = 10 x-10 = 1 + (y − 6)2, x = 10 - 10 x = 11 + (y − 6)2, x = 0 -x = 11 + (y − 6)2, -x = 0 x = -11 - (y − 6)2, x = 0

OpenStudy (anonymous):

Okay, what did you just do there ? lol

OpenStudy (anonymous):

i think it's better not to move it...

OpenStudy (amistre64):

changing x for y at the moment only orients it normally y = -11 - (x − 6)^2, y = 0 but yeah ... lets not try to move it this time :)

OpenStudy (amistre64):

define the function for height, and radius using the given graph as a guide

OpenStudy (anonymous):